OpenStudy (anonymous):

An elementary school is having a reading contest, to see which class can read the most number of minutes over the month. Each week, the teachers add up the number of hours their students have read at home and in school. What is the best way to display these weekly totals to show which class is ahead in the contest? A. histogram B. bar graph C. box-and-whisker plot D. ogive

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

any guess?

OpenStudy (anonymous):

A

OpenStudy (daniellelovee):

I dont think so because a histogram and a bar graph are pretty much the same thing.....

OpenStudy (anonymous):

so those things would be knocked out right?

OpenStudy (daniellelovee):

I guess so :P

OpenStudy (anonymous):

C?

OpenStudy (daniellelovee):

hold hold on I think you were right at first

OpenStudy (anonymous):

was I?

OpenStudy (daniellelovee):

yeah lol hold on Im thinking

OpenStudy (daniellelovee):

no nvm look http://www.mathsisfun.com/data/images/bar-chart-vs-histogram.gif

OpenStudy (daniellelovee):

notice the possibility?

OpenStudy (anonymous):

ok then my next guess is C

OpenStudy (daniellelovee):

-.-'

OpenStudy (daniellelovee):

lol you totally ignored the picture didnt you xD

OpenStudy (daniellelovee):

C and D dont have categories they just follow one path (meaning that you will only be able to see the results of one class)

OpenStudy (anonymous):

THEN ITS A!!! I'm positive. or B

OpenStudy (daniellelovee):

lol xD focus....

OpenStudy (daniellelovee):

LOOK AT THE PICTURE!!

OpenStudy (anonymous):

A..

OpenStudy (daniellelovee):

http://media1.giphy.com/media/mcMNAiDvn9tTO/200_s.gif

OpenStudy (anonymous):

B its B!

OpenStudy (daniellelovee):

lol according to the picture it is B because it has categories which we need to compare lol

OpenStudy (anonymous):

The median of the values in a data set is g. If 875 were subtracted from each of the values in the data set, what would be the median of the resulting data? A. g + 875 B. 875 − g C. g - 875 D. 875 ? g

OpenStudy (anonymous):

i got it right :D

OpenStudy (anonymous):

B is 875-g

OpenStudy (daniellelovee):

lol thankfully I do know my graphs :P :)

OpenStudy (daniellelovee):

the median does not change so this is tricky...

OpenStudy (daniellelovee):

Im guessing that you add it

OpenStudy (daniellelovee):

since fore example lets say we have 1,2,3,4,5 then 1+2+3+4+5=15/5=3 now lets say we add 2 to each value 3,4,5,6,7 3+4+5+6+7=25/5=5 so 5-2=2 which is the same increase as in the data

OpenStudy (daniellelovee):

@Tallan

OpenStudy (anonymous):

sorry internet went down

OpenStudy (anonymous):

i would say C @Daniellelovee

OpenStudy (anonymous):

Since you have whatever number it is and your subtracting from it

OpenStudy (daniellelovee):

I think you are correct :) based on what I explained up there lol and yeah is ok lol (not really I was like bruhh really?)

OpenStudy (anonymous):

The value of an explanatory variable is 8, while the corresponding value of the response variable is 18. What would be the coordinates of this data point when plotted on a scatterplot? A. (8, 18) B. (8, 26) C. (18, 8) D. (26, 8) I think A

OpenStudy (daniellelovee):

how many questions???

OpenStudy (daniellelovee):

and I agree with you

OpenStudy (anonymous):

theres 5 left. :P

OpenStudy (anonymous):

What is the first quartile of the data set represented by the box plot shown below?  A. 15 B. 20 C. 8 D. 35